The maximum amount of manganese(II) sulfide that will dissolve in a 0.186 M potassium sulfide solution is ___ M.

- The Ksp for manganese(II) sulfide is

5.1 × 10-15
Answer #1


K2S here is Strong electrolyte
It will dissociate completely to give [S2-] = 0.186 M
At equilibrium:
MgS       <---->     Mg2+     +          S2-    
                     s                  0.186 + s
Ksp = [Mg2+][S2-]
5.1*10^-15=(s)*(0.186+ s)
Since Ksp is small, s can be ignored as compared to 0.186
Above expression thus becomes:
5.1*10^-15=(s)*(0.186)
5.1*10^-15= (s) * 0.186
s = 2.742*10^-14 M
Answer: 2.7*10^-14 M